# VolleyBots: A Testbed for Multi-Drone Volleyball Game Combining Motion Control and Strategic Play

### v3

@@ Abstract
Robot sports, characterized by well-defined objectives, explicit rules, and dynamic interactions, present ideal scenarios for demonstrating embodied intelligence. In this paper, we present VolleyBots, a novel robot sports testbed where multiple drones cooperate and compete in the sport of volleyball under physical dynamics. VolleyBots integrates three features within a unified platform: competitive and cooperative gameplay, turn-based interaction structure, and agile 3D maneuvering. Competitive and cooperative gameplay challenges each drone to coordinate with its teammates while anticipating and countering opposing teams’ tactics. Turn-based interaction demands precise timing, accurate state prediction, and management of long-horizon temporal dependencies. Agile 3D maneuvering requires rapid accelerations, sharp turns, and precise 3D positioning despite the quadrotor’s underactuated dynamics. These intertwined features yield a complex problem combining motion control and strategic play, with no available expert demonstrations. We provide a comprehensive suite of tasks ranging from single-drone drills to multi-drone cooperative and competitive tasks, accompanied by baseline evaluations of representative multi-agent reinforcement learning (MARL) and game-theoretic algorithms. Simulation results show that on-policy reinforcement learning (RL) methods outperform off-policy methods in single-agent tasks, but both approaches struggle in complex tasks that combine motion control and strategic play. We additionally design a hierarchical policy which achieves 69.5% win rate against the strongest baseline in the 3 vs 3 task, underscoring its potential as an effective solution for tackling the complex interplay between low-level control and high-level strategy. The project page is at https://sites.google.com/view/thu-volleybots.
footnotetext: Equal contribution.footnotetext: Corresponding authors. {yuchao,yu-wang}@tsinghua.edu.cn
@@ 1 Introduction
The overview of the VolleyBots testbed is shown in Fig. 1. Built on Nvidia Isaac Sim [8], VolleyBots supports efficient GPU-based data collection. Inspired by how humans progressively learn the structure of volleyball, we design a curriculum of tasks ranging from single-drone drills to multi-drone cooperative plays and competitive matchups. We have also implemented multi-agent reinforcement learning (MARL) and game-theoretic baselines, and provided benchmark results. Simulation results show that on-policy reinforcement learning (RL) methods outperform off-policy counterparts in single-agent tasks. However, both approaches struggle in more complex tasks that require low-level motion control and high-level strategic play. We envision VolleyBots as a valuable platform for advancing the study of embodied intelligence in physically grounded, multi-agent robotic environments. Our main contributions are summarized as follows:
@@ 2 Related work
2.1 Robot sports
The integration of sensing, actuation, and autonomy has enabled a wide range of robotic platforms, spanning robotic arms, quadrupeds, humanoids, and aerial drones, to undertake increasingly complex tasks. Robot sports provide a compelling testbed for evaluating their capabilities within well-defined rule sets. A classic example is robot soccer: since the initiative of RoboCup [12], research into autonomous football has driven advances in multi-agent coordination, strategic planning, and hardware integration. Early approaches [13, 14] to robot sports relied primarily on classical control and planning techniques. With the growth of data-driven methods, imitation learning algorithms [15] enabled robots to learn complex motion policies directly from expert demonstrations. More recently, RL methods have also achieved remarkable performance. With RL, Researchers have explored a wide range of robot platforms for sports tasks. Robot arms on mobile bases have learned table tennis [5, 6] and badminton [9]. Quadrupeds have commanded basic soccer drills [2] and played in multi-agent football matches [4]. Humanoid robots have demonstrated competitive 1 vs 1 [3] and 2 vs 2 [1] football skills and participated in simulated Olympic-style events SMPLOlympics [10]. Drones have achieved human-surpassing racing performance [11] and tackled multi-UAV pursuit-evasion tasks with rule-based pursuit policies [7]. Despite these advances, there remains a need for environments that combine high-mobility platforms (e.g., drones) with mixed cooperative-competitive dynamics and require both high-level decision-making and low-level continuous control. To fill this gap, we introduce VolleyBots, a turn-based, drone-focused sports environment that seamlessly integrates strategic planning with agile control. Built on a realistic physics simulator, VolleyBots offers a unique testbed for advancing research in agile, decision-driven robot control. A detailed comparison between VolleyBots and representative learning-based robot sports platforms is provided in Table 1.
@@ 3.1 Simulation entity
[…]
@@ 5 Benchmark results
QMIX $0.09\pm 0.00$ $0.09\pm 0.00$ $0.02\pm 0.00$ $0.02\pm 0.00$ $0.02\pm 0.00$ $0.02\pm 0.00$
MADDPG $0.88\pm 0.07$ $0.90\pm 0.01$ $0.23\pm 0.01$ $0.24\pm 0.01$ $0.23\pm 0.00$ $0.23\pm 0.00$
MAPPO $\bm{11.17\pm 1.03}$ $\bm{14.05\pm 0.36}$ $\bm{0.25\pm 0.00}$ $\bm{0.99\pm 0.00}$ $\bm{0.25\pm 0.00}$ $0.75\pm 0.01$
HAPPO $6.65\pm 3.59$ $12.35\pm 0.67$ $\bm{0.25\pm 0.00}$ $0.98\pm 0.00$ $\bm{0.25\pm 0.00}$ $\bm{0.82\pm 0.11}$
MAT $7.68\pm 4.9$ $13.41\pm 0.14$ $\bm{0.25\pm 0.00}$ $\bm{0.99\pm 0.00}$ $\bm{0.25\pm 0.00}$ $0.75\pm 0.00$
@@ 5.1 Results of single-agent tasks
We evaluate five RL algorithms including Deep Q-Network (DQN) [27], Deterministic Policy Gradient (DDPG) [28], Twin Delayed DDPG (TD3) [29], Soft Actor-Critic (SAC) [30], and Proximal Policy Optimization (PPO) [31] in three single-agent tasks. We compare their performance under both CTBR and PRT action spaces. The averaged results over three seeds are shown in Table 2.
Using the same number of training frames, PPO vastly outperforms all other methods in every task under both action-space choices; by contrast, DQN fails entirely, and DDPG, TD3 and SAC achieve only moderate success. DQN fails because it’s limited to discrete actions, forcing coarse binning of continuous drone controls and losing precision. In contrast, while DDPG, TD3, and SAC can handle continuous actions, PPO’s clipped surrogate objective and on-policy updates provide greater stability and automatically adapt exploration, enabling it to outperform other single-agent methods.
Comparing different action spaces, the final results indicate that PRT slightly outperforms CTBR in most tasks. This outcome is likely due to PRT providing more granular control over each motor’s thrust, enabling the drone to maximize task-specific performance with precise adjustments. On the other hand, CTBR demonstrates a slightly faster learning speed in some tasks, as its higher-level abstraction simplifies the control process and reduces the learning complexity. For optimal task performance, we use PRT as the default action space in subsequent experiments. More experiment results and learning curves are provided in Appendix F.3.
@@ 5.2 Results of multi-agent cooperative tasks
We evaluate five MARL algorithms including QMIX [32], Multi-Agent DDPG (MADDPG) [33], Multi-Agent PPO (MAPPO) [34], Heterogeneous-Agent PPO (HAPPO) [35], Multi-Agent Transformer (MAT) [36] in three multi-agent cooperative tasks. We also compare their performance with and without reward shaping. The averaged results over three seeds are shown in Table 3.
@@ 5.3 Results of multi-agent competitive tasks
In the 1 vs 1 task, all algorithms manage to learn behaviors like reliably returning the ball and maintaining optimal positioning for subsequent volleys. However, the exploitability metric reveals that the learned policies are still far from achieving a Nash equilibrium, indicating limited robustness in this two-player zero-sum game. This performance gap highlights the inherent challenge of hierarchical decision-making in this task, where drones must simultaneously execute precise low-level motion control and engage in high-level strategic gameplay. This challenge presents new opportunities for designing algorithms that can better integrate hierarchical decision-making capabilities. In the 3 vs 3 task, algorithms learn to serve the ball but fail to develop more advanced team strategies. This outcome underscores the compounded challenges in this scenario, where each team of three drones needs to not only cooperate internally but also compete against the opposing team. The increased difficulty of achieving high-level strategic play in such a mixed cooperative-competitive environment further amplifies the hierarchical challenges observed in 1 vs 1.
@@ 6 Conclusion
In this work, we introduce VolleyBots, a novel multi-drone volleyball testbed that unifies competitive-cooperative gameplay, turn-based interaction, and agile 3D motion control within a high-fidelity physics simulation. Compounding these features, it demands both motion control and strategic play. Built atop NVIDIA Isaac Sim, VolleyBots offers a structured curriculum of tasks, from single-agent drills and multi-agent cooperative challenges to multi-agent competitive matches. To enable systematic benchmarking, VolleyBots provides implementations of both (MA)RL and game-theoretic baselines across these tasks. Our extensive benchmarks reveal that on-policy RL methods consistently outperform off-policy counterparts in low-level control tasks. However, both of them still struggle with the tasks demanding both mixed motion control and strategic play, especially in large-scale competitive matches. To address this, we design a simple hierarchical policy that decomposes strategy and control: in the 3 vs 3 task, it achieves a 69.5% win rate against the strongest baseline, highlighting the promise of hierarchical structures. Going forward, VolleyBots provides a challenging and versatile platform for advancing embodied intelligence in agile robotic systems, inviting novel algorithmic innovations that bridge motion control and strategic play in multi-agent domains.
@@
This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496.
@@ Appendix B Limitations
Despite the promising advances of VolleyBots in combining high-level strategic play with low-level motion control, our work has several limitations. First, all experiments were conducted in a high-fidelity simulator, and we have not evaluated the policy’s sim-to-real transferability, leaving its performance and robustness on physical UAV platforms untested. Second, we rely on fully state-based observations, which overlook challenges such as visual input. Finally, we did not include the traditional drone control algorithms which, despite their limitations in handling team-level strategy, aggressive maneuvers, and ball interactions, would provide useful baselines.
@@ C.1 Court
The volleyball court in our environment is depicted in Fig. 5. The court is divided into two equal halves by the $y$-axis, which serves as the dividing line separating the two teams. The coordinate origin is located at the midpoint of the dividing line, and the $x$-axis extends along the length of the court. The total court length is $18\,{m}$, with $x=-9$ and $x=9$ marking the ends of the court. The $y$-axis extends across the width of the court, with a total width of $9\,{m}$, spanning from $y=-4.5$ to $y=4.5$. The net is positioned at the center of the court along the $y$-axis, with a height of $2.43\,{m}$, and spans horizontally from $(0,-4.5)$ to $(0,4.5)$.
@@ C.2 Drone
We use the Iris quadrotor model [25] as the primary drone platform, augmented with a virtual “racket” of radius $0.2\,{m}$ and coefficient of restitution $0.8$ for ball striking. The drone’s root state is a vector with dimension $23$, including its position, rotation, linear velocity, angular velocity, forward orientation, upward orientation, and normalized rotor speeds.
@@ Appendix D Details of task design
Task Reward dist_to_target [0,0.5]×[0,0.5$\times$[ 0 , 0.5 ] × # step related to drone’s distance to the current target
[…]
@@
distance $[0,+\infty]$ related to the landing position’s distance to the anchor
dist_to_anchor $[-\infty,0]$ related to drone’s distance to the anchor
@@
[…]
Shaping Reward dist_to_ball_xy [0,1]×[0,1$\times$[ 0 , 1 ] × # step related to drone’s horizontal distance to the ball
dist_to_ball_z [0,1]×[0,1$\times$[ 0 , 1 ] × # step related to drone’s clipped vertical distance to the ball
@@
[…]
dist_to_anchor $[-\infty,0]$ related to drone’ distance to its anchor
[…]
dist_to_ball [0,0.05]×[0,0.05$\times$[ 0 , 0.05 ] × # step related to drone’s distance to the ball
@@
[…]
@@
dist_to_anchor $[-\infty,0]$ related to drone’s distance its anchor
[…]
spike_velociy [0,+]×[0,$\infty$]$\times$[ 0 , + ∞ ] × # spike related to ball’s downward velocity after spike
dist_to_ball [0,0.05]×[0,0.05$\times$[ 0 , 0.05 ] × # step related to drone’s distance to the ball
@@
[…]
@@
dist_to_anchor $[-\infty,0]$ related to drone’s distance to its anchor
[…]
spike_velociy [0,+]×[0,$\infty$]$\times$[ 0 , + ∞ ] × # spike related to ball’s downward velocity after spike
dist_to_ball [0,0.05]×[0,0.05$\times$[ 0 , 0.05 ] × # step related to drone’s distance to the ball
@@
drone_out_of_court [0,0.2]×[0,0.2$\times$[ 0 , 0.2 ] × # step related to drone’s distance out of its court
@@
[…]
dist_to_ball [0,0.5]×[0,0.5$\times$[ 0 , 0.5 ] × # step related to drone’s distance to the ball
@@
Shaping Reward success_hit $0\}\times$ , 10 } × # hit drone hits the ball
dist_to_anchor [0,0.05]×[0,0.05$\times$[ 0 , 0.05 ] × # step related to drone’s distance to its anchor
dist_to_ball [0,0.5]×[0,0.5$\times$[ 0 , 0.5 ] × # step related to drone’s distance to the ball
@@
Shaping Reward success_hit $0\}\times$ , 10 } × # hit drone hits the ball
dist_to_anchor [0,0.05]×[0,0.05$\times$[ 0 , 0.05 ] × # step related to drone’s distance to its anchor
dist_to_ball [0,0.5]×[0,0.5$\times$[ 0 , 0.5 ] × # step related to drone’s distance to the ball

### v6

@@ Abstract
Robot sports, characterized by well-defined objectives, explicit rules, and dynamic interactions, present ideal scenarios for demonstrating embodied intelligence. In this paper, we present VolleyBots, a novel robot sports testbed where multiple drones cooperate and compete in the sport of volleyball under physical dynamics. VolleyBots integrates three features within a unified platform: competitive and cooperative gameplay, turn-based interaction structure, and agile 3D maneuvering. These intertwined features yield a complex problem combining motion control and strategic play, with no available expert demonstrations. We provide a comprehensive suite of tasks ranging from single-drone drills to multi-drone cooperative and competitive tasks, accompanied by baseline evaluations of representative reinforcement learning (RL), multi-agent reinforcement learning (MARL) and game-theoretic algorithms. Simulation results show that on-policy RL methods outperform off-policy methods in single-agent tasks, but both approaches struggle in complex tasks that combine motion control and strategic play. We additionally design a hierarchical policy which achieves 69.5% win rate against the strongest baseline in the 3 vs 3 task, demonstrating its potential for tackling the complex interplay between low-level control and high-level strategy. To highlight VolleyBots’ sim-to-real potential, we further demonstrate the zero-shot deployment of a policy trained entirely in simulation on real-world drones.
Benchmark & Code: https://github.com/thu-uav/VolleyBots
Project Website: https://volleybots.github.io
footnotetext: Equal contribution. {zelai.eecs,jimmyzhangruize}@gmail.comfootnotetext: Corresponding authors. zoeyuchao@gmail.com, yu-wang@mail.tsinghua.edu.cn
@@ 1 Introduction
The overview of the VolleyBots testbed is shown in Fig. 1. Built on Nvidia Isaac Sim [26], VolleyBots supports efficient GPU-based data collection. Inspired by how humans progressively learn the structure of volleyball, we design a curriculum of tasks ranging from single-drone drills to multi-drone cooperative plays and competitive matchups. We have also implemented reinforcement learning (RL), multi-agent reinforcement learning (MARL) and game-theoretic baselines, and provided benchmark results. In single-agent tasks, simulation results show that with a single set of hyperparameters, on-policy RL methods maintains consistently strong performance across multiple tasks, demonstrating superior robustness compared to off-policy methods. However, both approaches struggle in more complex tasks that require low-level motion control and high-level strategic play. To demonstrate real-world deployment ability, we show a policy trained to bump volleyball can be deployed on an open-source quadrotor equipped with a racket in a zero-shot manner. We envision VolleyBots as a valuable platform for advancing the study of embodied intelligence in physically grounded, multi-agent robotic environments.
Our main contributions are summarized as follows:
@@ 2 Related work
2.1 Robot sports
The integration of sensing, actuation, and autonomy has enabled a wide range of robotic platforms, spanning robotic arms, quadrupeds, humanoids, and aerial drones, to undertake increasingly complex tasks. Robot sports provide a compelling testbed for evaluating their capabilities within well-defined rule sets. A classic example is robot soccer: since the initiative of RoboCup [17], research into autonomous football has driven advances in multi-agent coordination, strategic planning, and hardware integration. Early approaches [1, 28] to robot sports relied primarily on classical control and planning techniques. With the growth of data-driven methods, imitation learning algorithms [29] enabled robots to learn complex motion policies directly from expert demonstrations. More recently, RL methods have also achieved remarkable performance. With RL, Researchers have explored a wide range of robot platforms for sports tasks. Robot arms on mobile bases have learned table tennis [6, 24] and badminton [36]. Quadrupeds have commanded basic soccer drills [14] and played in multi‐agent football matches [40]. Humanoid robots have demonstrated competitive 1 vs 1 [10] and 2 vs 2 [21] football skills and participated in simulated Olympic‐style events SMPLOlympics [23]. Drones have achieved human‐surpassing racing performance [16] and tackled multi-UAV pursuit-evasion tasks with rule-based pursuit policies [3]. Despite these advances, there remains a need for environments that combine high-mobility platforms (e.g., drones) with mixed cooperative-competitive dynamics and require both high-level decision-making and low-level continuous control. To fill this gap, we introduce VolleyBots, a turn-based, drone-focused sports environment that seamlessly integrates strategic planning with agile control. Built on a realistic physics simulator, VolleyBots offers a unique testbed for advancing research in agile, decision-driven robot control. A detailed comparison between VolleyBots and representative learning-based robot sports platforms is provided in Table 1.
@@ 3.1 Simulation entity
Our environment simulates real-world physics dynamics and interactions of three key components including the drones, the ball, and the court. We provide a flexible configuration of each entity’s model and parameters to enable a wide range of task designs. For the default configuration, we adopt the Iris quadrotor model [9] as the primary drone platform, augmented with a virtual “racket” of radius $0.2\,\text{m}$ and coefficient of restitution $0.8$ for ball striking. The ball is modeled as a sphere with a radius of $0.1\,\text{m}$, a mass of $5\,\text{g}$, and a coefficient of restitution of $0.8$, enabling realistic bounces and interactions with both drones and the environment. The court follows standard volleyball dimensions of $9\,\text{m}\times 18\,\text{m}$ with a net height of $2.43\,\text{m}$.
@@ 5 Benchmark results
QMIX $09\pm 0.01 0.0$$9\pm 0.00 0.0$$2\pm 0.00 0.0$$2\pm 0.00 0.0$$2\pm 0.00 0.0$$02\pm 0.00$
MADDPG $0.79\pm 0.15$ $0.84\pm 0.09$ $0.22\pm 0.02$ $3\pm 0.01 0.22±0$$2\pm 0.02$ $0.22\pm 0.02$
MAPPO $\bm{11.32\pm 0.91}$ $\bm{13.71\pm 0.58}$ $\bm{0.25\pm 0.00}$ $\bm{0.99\pm 0.00}$ $\bm{0.25\pm 0.00}$ $0.75\pm 0.01$
HAPPO $7.95\pm 3.67$ $12.14\pm 0.83$ $\bm{0.25\pm 0.00}$ $0.98\pm 0.00$ $\bm{0.25\pm 0.00}$ $0.79\pm 0.10$
MAT $7.39\pm 6.00$ $13.11\pm 0.43$ $\bm{0.25\pm 0.00}$ $0.89\pm 0.13$ $\bm{0.25\pm 0.00}$ $\bm{0.80\pm 0.11}$
@@ 5.1 Results of single-agent tasks
We evaluate five RL algorithms including Deep Q-Network (DQN) [27], Deterministic Policy Gradient (DDPG) [20], Twin Delayed DDPG (TD3) [8], Soft Actor-Critic (SAC) [11], and Proximal Policy Optimization (PPO) [33] in three single-agent tasks. We compare their performance under both CTBR and PRT action spaces. The averaged results over five seeds are shown in Table 2.
For each algorithm, the same set of hyperparameters is used across all tasks to assess its cross-task robustness, while different algorithms are independently tuned for fairness. Details of the hyperparameter tuning process are provided in Appendix F.2.1. Under this setup, PPO consistently outperforms all other methods in every task and under both action-space configurations; by contrast, DQN fails entirely, and DDPG, TD3 and SAC achieve only moderate success. DQN fails because it’s limited to discrete actions, forcing coarse binning of continuous drone controls and losing precision. In contrast, while DDPG, TD3, and SAC can handle continuous actions, PPO’s clipped surrogate objective and on-policy updates provide greater stability and adaptive exploration, leading to superior performance and stronger cross-task robustness to hyperparameter settings.
Comparing different action spaces, the final results indicate that PRT slightly outperforms CTBR in most tasks. This outcome is likely due to PRT providing more granular control over each motor’s thrust, enabling the drone to maximize task-specific performance with precise adjustments. On the other hand, CTBR demonstrates a slightly faster learning speed in some tasks, as its higher-level abstraction simplifies the control process and reduces the learning complexity. For optimal task performance, we use PRT as the default action space in subsequent experiments. Additional experimental results and learning curves are presented in Appendix F.3.
@@ 5.2 Results of multi-agent cooperative tasks
We evaluate five MARL algorithms including QMIX [31], Multi-Agent DDPG (MADDPG) [22], Multi-Agent PPO (MAPPO) [43], Heterogeneous-Agent PPO (HAPPO) [18], Multi-Agent Transformer (MAT) [38] in three multi-agent cooperative tasks. We also compare their performance with and without reward shaping. The averaged results over five seeds are shown in Table 3.
@@ 5.3 Results of multi-agent competitive tasks
In the 1 vs 1 task, all algorithms manage to learn behaviors like reliably returning the ball and maintaining optimal positioning for subsequent volleys. However, the exploitability metric reveals that the learned policies are still far from achieving a Nash equilibrium, indicating limited robustness in this two-player zero-sum game. This performance gap highlights the inherent challenge of hierarchical decision-making in this task, where drones must simultaneously execute precise low-level motion control and engage in high-level strategic gameplay. This challenge presents new opportunities for designing algorithms that can better integrate hierarchical decision-making capabilities. In the 3 vs 3 task, algorithms learn to serve the ball but fail to develop more advanced team strategies. This outcome underscores the compounded challenges in this scenario, where each team of three drones needs to not only cooperate internally but also compete against the opposing team. The increased difficulty of achieving high-level strategic play in such a mixed cooperative-competitive environment further amplifies the hierarchical challenges observed in the 1 vs 1 task.
@@ 5.4 Hierarchical policy
6 Sim-to-Real
Figure 5: Zero-shot sim-to-real experiment on the Solo Bump task.
We use the Solo Bump task as a demonstration of the policy’s ability to zero-shot transfer to the real world. We use a quadrotor with a rigidly mounted badminton racket. The state of both the drone and the ball is captured using a motion capture system. The drone is modeled as a rigid body, with its position and orientation provided by the motion capture system. The drone’s velocity is estimated using an Extended Kalman Filter (EKF) that fuses pose data from the motion capture system and IMU data from the PX4 Autopilot. The ball is modeled as a point mass, with its position sent by the motion capture system and its velocity indirectly computed through a Kalman Filter. The drone’s dynamics parameters and the ball’s properties are determined through system identification. To simulate real-world noise and imperfect execution of actions, small randomizations are introduced in the ball’s initial position, coefficient of restitution, and the ball’s rebound velocity after each collision with the drone. Inspired by [4], we also add a smoothness reward to encourage smooth actions. The policy uses CTBR as output and is deployed on the onboard Nvidia Orin processor. As shown in Fig. 5, experiment results show that the drone successfully performs bump tasks multiple times, providing initial evidence of sim-to-real transfer capability. The real-world deployment videos are publicly available on our project website.
@@ 7 Conclusion
In this work, we introduce VolleyBots, a novel multi‐drone volleyball testbed that unifies competitive-cooperative gameplay, turn-based interaction, and agile 3D motion control within a high-fidelity physics simulation. Compounding these features, it demands both motion control and strategic play. Built atop NVIDIA Isaac Sim, VolleyBots offers a structured curriculum of tasks, from single-agent drills and multi-agent cooperative challenges to multi-agent competitive matches. To enable systematic benchmarking, VolleyBots provides implementations of both (MA)RL and game-theoretic baselines across these tasks. Our extensive benchmarks reveal that on-policy RL methods consistently outperform their off-policy counterparts in low-level control tasks, and exhibit stronger cross-task robustness under a single set of hyperparameters. However, both of them still struggle with the tasks demanding both mixed motion control and strategic play, especially in large-scale competitive matches. To address this, we design a simple hierarchical policy that decomposes strategy and control: in the 3 vs 3 task, it achieves a 69.5% win rate against the strongest baseline, highlighting the promise of hierarchical structures. We also showcase the feasibility of deploying policies trained in simulations directly onto physical drones, emphasizing VolleyBots’ sim-to-real transfer and practical utility in real-world applications. Going forward, VolleyBots provides a challenging and versatile platform for advancing embodied intelligence in agile robotic systems, inviting novel algorithmic innovations that bridge motion control and strategic play in multi-agent domains.
@@
This research was supported by National Natural Science Foundation of China (No.62406159, 62325405), Postdoctoral Fellowship Program of CPSF under Grant Number (GZC20240830, 2024M761676), China Postdoctoral Science Special Foundation 2024T170496. Additional support was provided by Tsinghua-Efort Joint Research Center for EAI Computation and Perception, Beijing National Research Center for Information Science and Technology (BNRist), Beijing Innovation Center for Future Chips, and State Key laboratory of Space Network and Communications.
@@ Appendix B Limitations
Despite the promising advances of VolleyBots in combining high-level strategic play with low‐level motion control, our work has several limitations. First, apart from the Solo Bump task, the sim-to-real transferability of the learned policies in other tasks has not yet been evaluated on physical UAV platforms. Second, we rely on fully state-based observations, which overlook challenges such as visual input. Finally, traditional drone control algorithms were not included. Although they struggle with team-level coordination, aggressive maneuvers, and ball interactions, they could still provide informative baselines.
@@ C.1 Court
The volleyball court in our environment is depicted in Fig. 6. The court is divided into two equal halves by the $y$-axis, which serves as the dividing line separating the two teams. The coordinate origin is located at the midpoint of the dividing line, and the $x$-axis extends along the length of the court. The total court length is 18m18\,{m}, with x=9x=-9 and $x=9$ marking the ends of the court. The $y$-axis extends across the width of the court, with a total width of 9m9\,{m}, spanning from y=4.5y=-4.5 to $y=4.5$. The net is positioned at the center of the court along the $y$-axis, with a height of 2.43m2.43\,{m}, and spans horizontally from (0,4.5)(0,-4.5) to $(0,4.5)$.
@@ C.2 Drone
We use the Iris quadrotor model [9] as the primary drone platform, augmented with a virtual “racket” of radius 0.2m0.2\,{m} and coefficient of restitution $0.8$ for ball striking. The drone’s root state is a vector with dimension $23$, including its position, rotation, linear velocity, angular velocity, forward orientation, upward orientation, and normalized rotor speeds.
@@ Appendix D Details of task design
Task Reward dist_to_target $0,0.5]\times$ # step related to drone’s distance to the current target
target_stay $\{0,2.5\}\times$ # in_target drone stays in the current target region
@@
distance $0,+\infty]$ related to the landing position’s distance to the anchor
dist_to_anchor $\infty,0]$ related to drone’s distance to the anchor
@@
Task Reward success_hit $\{0,1\}\times$ # hit drone hits the ball
success_height $\{0,8\}\times$ # hit ball reaches the minimum height
Shaping Reward dist_to_ball_xy $0,1]\times$ # step related to drone’s horizontal distance to the ball
dist_to_ball_z $0,1]\times$ # step related to drone’s clipped vertical distance to the ball
@@
Task Reward success_hit $\{0,1\}\times$ # hit drone hits the ball
success_cross $\{0,1\}\times$ # hit ball crosses the height
dist_to_anchor $\infty,0]$ related to drone’ distance to its anchor
Shaping Reward hit_direction $\{0,1\}\times$ # hit drone hits the ball towards the other drone
dist_to_ball $0,0.05]\times$ # step related to drone’s distance to the ball
@@
Task Reward success_hit $\{0,5\}\times$ # hit drone hits the ball
downward_spike $\{0,5\}\times$ # spike ball’s velocity is downward after spike
@@
dist_to_anchor $\infty,0]$ related to drone’s distance its anchor
Shaping Reward hit_direction $\{0,1\}\times$ # hit drone hits the ball towards its target
spike_velociy $0,+\infty]\times$ # spike related to ball’s downward velocity after spike
dist_to_ball $0,0.05]\times$ # step related to drone’s distance to the ball
@@
Task Reward success_hit $\{0,5\}\times$ # hit drone hits the ball
downward_spike $\{0,5\}\times$ # spike ball’s velocity is downward after spike
@@
dist_to_anchor $\infty,0]$ related to drone’s distance to its anchor
Shaping Reward hit_direction $\{0,1\}\times$ # hit drone hits the ball towards their targets
spike_velociy $0,+\infty]\times$ # spike related to ball’s downward velocity after spike
dist_to_ball $0,0.05]\times$ # step related to drone’s distance to the ball
@@
drone_out_of_court $0,0.2]\times$ # step related to drone’s distance out of its court
@@
Shaping Reward success_hit $\{0,5\}\times$ # hit drone hits the ball
dist_to_ball $0,0.5]\times$ # step related to drone’s distance to the ball
@@
Shaping Reward success_hit $\{0,10\}\times$ # hit drone hits the ball
dist_to_anchor $0,0.05]\times$ # step related to drone’s distance to its anchor
dist_to_ball $0,0.5]\times$ # step related to drone’s distance to the ball
@@
Shaping Reward success_hit $\{0,10\}\times$ # hit drone hits the ball
dist_to_anchor $0,0.05]\times$ # step related to drone’s distance to its anchor
dist_to_ball $0,0.5]\times$ # step related to drone’s distance to the ball